OpenStudy (anonymous):

HELP ! If the dimensions of a rectangle become one-fifth of the original dimensions, then area of the new rectangle will be _______of the area of the original rectangle.

OpenStudy (paxpolaris):

let's take an example: all sides of the original rectangle are 1 in. (so, technically it's a square)

OpenStudy (paxpolaris):

so the area of the original would be = 1in * 1 in = 1 sq. in.

OpenStudy (anonymous):

so now would we do 1 * 1/5 = __

OpenStudy (anonymous):

so .2? is that one

OpenStudy (paxpolaris):

in the new rectangle each side would be 1/5 so new area = \[\frac 15 \times \frac 15 = \Large{1 \over 25}\]

OpenStudy (anonymous):

so the area of the new rectangle will be 1/25th of the area of the original rectangle?

OpenStudy (paxpolaris):

yes .. if you are scaling something the ratio of the areas will be the square of the ratio of the sides... (i.e. the square of the scale factor)

OpenStudy (anonymous):

Oh! makes sense now! Thank you!